A stock solution of CuCl2 was made by adding 510g of solute to 0.95L of water. What volume of stock solution would you use to create 14.1L of 0.58M CuCl2 solution? Round your answer to the nearest 0.1L.
1
Expert's answer
2020-05-12T11:34:36-0400

n1(CuCl2)=m(CuCl2)/Mr(CuCl2)=510/135=3,78 moll

C1=n1/v1=3,78/0,95=3,98 moll/L

C1*V1=C2*V2

3,98*x=0,58*14,1

x=2,05477=> 2,1 liter